OpenStudy (anonymous):

x² - 6x + 4 = 0 solve and show work.

OpenStudy (anonymous):

x² - 6x + 4 = 0 x² - 6x = -4 x² - 6x + 9 = -4 +9 (x-3)^2 = 5 x-3 = +\[\sqrt{5}\] or x-3 = -\[\sqrt{5}\] x = 3+sqrt(5) or x = 3- sqrt(5)